How do you solve this?
Directions:
Write an equation of the line that passes through the given point and is parallel to the given line.
Questions:
1. (1,7), -6x + y = -1
2. (-2,5), 2y = 4x - 6
I would really appreciate it if someone helped me by doing a step-by-step process, I was really tired and unfortunately wasn't paying attention in class when we were learning this, thank you.

Answer by Alan3354(69443)

Directions:
Write an equation of the line that passes through the given point and is parallel to the given line.
Questions:
1. (1,7), -6x + y = -1
Find the slope of the line, m.
Rewrite it into slope-intercept form, that means solve for y.
-6x + y = -1
y = 6x - 1
The slope is the coefficient of x.
m = 6
Parallel lines have the same slope.
----
Use y-y1 = m*(x-x1) where (x1,y1) is the point.
y - 7 = 6*(x-1)
=========================
2. (-2,5), 2y = 4x - 6
Same as #1.